on a map,1 centimeter representsn80 kilometers.how many centimeters represent 280 kilometers
Answer by stanbon(75887)

on a map,1 centimeter represents 80 kilometers.
how many centimeters represent 280 kilometers
-------------------------------
x/280 = 1/80
----
x = 280/80
x = 3.5 cm
================
Cheers,
Stan H.